# ANSR-DT: An Adaptive Neuro-Symbolic Learning and Reasoning Framework for Digital Twins thanks: This research was partially supported by the U.S. National Science Foundation through Grant Nos. 2317117 and 2309760. (Corresponding author: Safayat Bin Hakim.)

### v2

@@ III-F Operational Example: Key Event Detection
In this operational scenario, the framework detected a pressure reading of 18.0, which activated neural_rule_1 and neural_rule_4 through the symbolic reasoning layer. The system identified this as a potential deviation with 99.81% confidence. Extended PPO training allowed the system to refine its control response from conservative adjustments to more decisive actions, demonstrating the value of reinforcement learning in industrial control scenarios.
This example demonstrates how ANSR-DT integrates neural time-series pattern extraction, symbolic reasoning, and reinforcement learning to detect operational deviations, make interpretable decisions, and adapt to changing conditions in real-time.
@@ IV Implementation
This section details the technical realization of ANSR-DT, emphasizing our methodologies for synthetic data generation, component development, and system integration. The complete implementation, including documentation and example scripts, is available as open-source at: https://github.com/sbhakim/ansr-dt.
@@ IV-A Synthetic Data Generation and Rationale
To evaluate ANSR-DT in controlled yet realistic conditions, we developed a synthetic data generation module that simulates industrial manufacturing scenarios. This approach offers several advantages over real-world data for initial development, including perfect ground truth labeling, systematic testing capabilities, and reproducibility for comparative evaluation. Our synthetic environment allows us to validate system performance without the privacy and security concerns associated with sensitive industrial data.
The updated dataset comprises 5,000 samples at 5-minute intervals, featuring seven variables: temperature, vibration, pressure, operational_hours (system uptime in hours), efficiency_index (normalized deviation from baseline), system_state (categorical operational mode), and performance_score (inverse efficiency). Key events, injected at a 5% rate, mimic operational thresholds (e.g., pressure spikes), maintenance events (e.g., vibration reductions), and state transitions (e.g., pressure shifts), generated via a multivariate normal distribution with correlations (0.3–0.5) and smoothed with Savitzky-Golay filters [24]. This design balances realism with computational efficiency, prioritizing rapid prototyping while new features like operational_hours and system_state enhance contextual modeling for RL and symbolic reasoning.
For validation, we conducted statistical tests including KS test and Anderson-Darling test [28], supplemented by cross-correlation analysis to ensure data fidelity. The visual representations in Fig. 3 illustrate the characteristics of our synthetic data, including sensor value distributions, time-series patterns, and dynamic changes.
@@ IV-B1 Deep Learning Component
The CNN-LSTM comprises two Conv1D layers with 64 and 128 filters (kernel size 3), followed by batch normalization and max-pooling, then two LSTM layers with 100 and 50 units, respectively, and an attention mechanism to prioritize critical time steps. This configuration totals 189,481 parameters, balancing complexity and efficiency for the 5,000-sample dataset.
For optimization, we employed Bayesian optimization [29] to tune hyperparameters including learning rates (1e-4 to 1e-2), dropout rates (0.1 to 0.5), and network architecture configurations. The final configuration used a learning rate of 1e-3 and dropout rate of 0.3. We utilized time series cross-validation with a walk-forward approach to ensure robust performance on sequential data, implementing early stopping (patience=10 epochs) to prevent overfitting. The model was trained using the Adam optimizer with batch size 64 for 20 epochs.
@@ IV-B2 Symbolic Reasoning Component
The symbolic component translates neural network outputs into interpretable logical rules and facts using a Prolog-compatible format [26]. Our knowledge base contains domain-specific predicates such as sensor_reading(Sensor, Value, Time), system_state(State, Time), and key_event(Type, Confidence, Time) to represent system state and observations.
The framework extracted 14 rules from test predictions, with the rule count stabilizing at 14 across PPO training durations of 10,240 and 200,704 timesteps, showing robustness. This consistency, despite varying RL policies, highlights the symbolic component’s reliability in translating neural predictions into interpretable rules on a condensed dataset.
@@
For example, neural_rule_1 flags operational deviations (confidence 0.999) when all metrics—temperature, vibration, pressure, efficiency_index—fall below their thresholds, as activated in the inference scenario (pressure=18.0), demonstrating actionable symbolic reasoning, while neural_rule_4 focuses on pressure-specific conditions:
For each candidate rule, we calculated confidence scores using the formula:
@@
where support(r) is the frequency of rule r’s condition being satisfied, support(r $\Rightarrow$ c) is the frequency of both the condition and conclusion being satisfied, and TP/(TP+FP) is the precision of the rule’s predictions. Rules with confidence scores above threshold $\tau$ = 0.85 were incorporated into the knowledge base, with conflicting rules resolved through confidence-based prioritization and specificity analysis. Most of the learned rules achieved confidence scores above 0.9, demonstrating high reliability in rule extraction.
@@ IV-B3 Neuro-Symbolic Integration Process
The integration between neural and symbolic components is implemented through a bidirectional interface that ensures seamless information flow [17]. Neural network outputs are converted into symbolic facts using a thresholding mechanism, where predictions exceeding a specified confidence threshold become symbolic facts for reasoning. Patterns detected by the neural component are analyzed to extract potential rules, which are evaluated for confidence against the observed data. Rules exceeding the confidence threshold $\tau$ are added to the knowledge base.
Symbolic insights remained identical across PPO training runs of 10,240 and 200,704 timesteps, consistently including System Stress Triggered and activations of neural_rule_1 and neural_rule_4. This stability, despite a refined RL policy, enhances interpretability by ensuring reliable, repeatable explanations for human operators, a key advantage in safety-critical settings.
@@ IV-C Reinforcement Learning Implementation
ANSR-DT implements the PPO algorithm [8] for adaptive decision-making in a custom OpenAI Gym-compatible environment. The environment encapsulates sensor readings, system states, and derived metrics as observation space, while providing operational adjustments as action space. Each state combines raw sensor readings, processed features, and symbolic facts to provide a comprehensive view of the system. The action space includes both continuous control parameters (e.g., operational settings, thresholds) and discrete decisions (e.g., maintenance scheduling, mode switching).
The multi-objective reward function balances efficiency, user satisfaction, and safety:
@@
where $\alpha_{1}=0.5$, $\alpha_{2}=0.3$, and $\alpha_{3}=0.2$ weight the relative importance of each objective. The policy network is implemented as a multi-layer perceptron with three hidden layers (256, 128, 64 units) that maps states to action probabilities. Our training configuration utilized a clipping parameter $\epsilon$ = 0.2, discount factor $\gamma$ = 0.99, GAE parameter $\lambda$ = 0.95, and learning rate 1e-3, with 10 epochs per policy update and batch size 64.
@@ IV-D Dynamic Rule Updating Mechanism
We implemented a continuous rule updating mechanism to adapt the symbolic knowledge base based on reinforcement learning insights. This process begins with rule candidate generation, where the neural component and RL agent experiences produce potential new rules. Each candidate rule’s confidence is assessed based on its predictive accuracy and support in the observed data. Rules exceeding the confidence threshold are incorporated into the knowledge base:
@@
Conflicting rules are resolved through confidence-based prioritization and specificity analysis, while a periodic rule pruning process removes rules with decreasing confidence or relevance to maintain knowledge base efficiency. This dynamic updating ensures that the symbolic reasoning component evolves alongside changing operational conditions and learning insights.
@@
IV-E Application Scenarios
To demonstrate the practical utility of ANSR-DT, we implemented three industrial application scenarios derived from our synthetic data. The first scenario focuses on predictive maintenance, where ANSR-DT monitors equipment health through temperature and vibration sensors to predict maintenance needs before critical failures. The neural component identifies vibration signatures indicative of bearing wear, while symbolic reasoning provides interpretable maintenance recommendations using rules such as: high_frequency_vibration(X) $\land$ rising_temperature(X) $\rightarrow$ maintenance_required(X, urgent). The system learns optimal maintenance scheduling based on operational constraints and historical outcomes.
The second scenario addresses process optimization, focusing on optimizing operational parameters for efficiency. This implementation integrates temperature, pressure, and derived efficiency metrics to monitor process performance. The RL component learns optimal operational settings under varying conditions, while symbolic rules explain why specific settings are optimal in given contexts. The system adapts to operator preferences while maintaining safety constraints.
The third scenario demonstrates deviation detection and response capabilities [30]. This implementation uses multi-sensor fusion to analyze temperature, vibration, and pressure data for complex event patterns. The system provides graduated responses prioritized by event severity and confidence, continuously learning to adapt to new deviation patterns. Symbolic reasoning offers operators clear explanations of detected operational deviations and recommended responses.
IV-F Integration and System Architecture
We integrated all components into a cohesive pipeline with well-defined interfaces between layers, as illustrated in Fig. 4. The implementation includes a data preprocessing pipeline for sensor data normalization and fusion, the neural component implemented in TensorFlow 2.8, a symbolic reasoning engine using ProbLog 2.2 [26], and a reinforcement learning module based on Stable Baselines3. These components are connected through a Python-based middleware with standardized interfaces, supplemented by visualization tools for monitoring system performance and explaining decisions.
@@ IV-G Future Extensions
Based on our implementation experience, we identified several promising directions for future enhancements. Future efforts should scale to larger datasets to test rule management beyond 14 rules and diversify rule extraction to capture more complex patterns, leveraging extended PPO training for enhanced adaptability. Additional extensions include adapting the framework for deployment on edge devices with constrained computational resources, integrating natural language processing for more intuitive human-machine interaction, and evaluating the framework’s applicability in diverse domains beyond manufacturing [31]. These extensions aim to further enhance the adaptability, interpretability, and practical utility of the ANSR-DT framework for complex industrial applications.
@@ V-A Experimental Setup
Experiments were conducted using synthetic industrial data with a 60/20/20 train/validation/test split containing 5% labeled dynamic patterns distributed across operational thresholds, maintenance patterns, state transitions, and interaction deviations. To address class imbalance in the dataset, we applied class weighting (0: 64.7, 1: 0.5) during model training. We established a rigorous comparison framework against multiple baseline models.
@@
Accuracy (%) 99.5 (peak) 86.87 +12.63 $<$ 0.01
Precision 0.80 $\pm$ 0.05 0.9358 -14.50 $<$ 0.05
Recall 0.75 $\pm$ 0.04 0.6355 +17.82 $<$ 0.05
@@
V-B Performance Analysis and Time-Series Pattern Extraction
Table II presents the comparative performance metrics between ANSR-DT and the CNN-LSTM baseline. Our framework achieves a peak validation accuracy of 99.5% by epoch 15 across 20 training epochs, compared to 86.87% for the CNN-LSTM baseline. However, fluctuations (e.g., dipping to 55% at epoch 16) suggest potential overfitting on the reduced dataset, mitigated by early stopping and class weighting to address imbalance.
The most substantial improvement appears in recall performance, with a 17.82% increase (p $<$ 0.05), indicating significantly enhanced capability to identify critical patterns with fewer false negatives—a crucial consideration for industrial safety applications.
This high detection accuracy is achieved through the combination of CNN-LSTM architecture with attention mechanism, which effectively learns temporal dependencies across multiple sensor streams. The precision-recall trade-off (shown in Fig. LABEL:fig:precision_recall) was deliberately engineered to prioritize recall in safety-critical industrial settings, where false negatives typically carry higher operational risk than false positives.
Figure 6 provides visual confirmation of classification performance. The confusion matrix (Fig. LABEL:fig:confusion_matrix) shows high true positive and true negative rates, while the precision-recall curve (Fig. LABEL:fig:precision_recall) demonstrates strong performance across different decision thresholds with an Average Precision score of 1.00. The training progression in Figure 5 shows convergence with validation accuracy reaching a peak of 99.5% at epoch 15, indicating efficient learning with some fluctuations due to the smaller dataset size.

### v3

@@ III-F Operational Example: Key Event Detection
An operational scenario illustrates ANSR-DT’s integrated decision-making process. The framework detected a pressure reading of 18.0 kPa, which activated neural_rule_1 and neural_rule_4 through the symbolic reasoning layer. The system identified this measurement as a potential operational deviation with 99.81% confidence based on the combined evidence from multiple sensor streams and activated symbolic rules. Extended PPO training allowed the system to refine its control response from conservative adjustments (small temperature modifications, minimal vibration changes) to more decisive actions (larger parameter adjustments aimed at rapidly restoring nominal operating conditions), demonstrating the value of reinforcement learning in industrial control scenarios. This example shows how ANSR-DT integrates neural time-series pattern extraction, symbolic reasoning, and reinforcement learning to detect operational deviations, make interpretable decisions, and adapt control strategies to changing conditions in real-time.
@@ IV Implementation
This section details the technical realization of ANSR-DT, emphasizing methodologies for synthetic data generation, component development, and system integration. The complete implementation, including documentation and example scripts, is available as open-source at https://github.com/sbhakim/ansr-dt.
@@ IV-A Synthetic Data Generation and Rationale
To evaluate ANSR-DT in controlled yet realistic conditions, we developed a synthetic data generation module that simulates industrial manufacturing scenarios. This approach offers perfect ground truth labeling, systematic testing capabilities, and reproducibility for comparative evaluation. The synthetic environment allows validation of system performance without the privacy and security concerns associated with sensitive industrial data, while enabling precise control over the types and frequencies of operational deviations injected for testing.
The dataset comprises 5,000 samples collected at 5-minute intervals, featuring seven variables that capture key aspects of industrial operation. Temperature, vibration, and pressure measurements represent direct sensor readings from physical equipment. Operational hours track system uptime to capture aging and wear effects. The efficiency index quantifies normalized deviation from baseline performance. System state encodes categorical operational modes (startup, normal operation, maintenance, shutdown). Performance score represents inverse efficiency as a derived metric for optimization objectives. Key events, injected at a 5% rate, mimic operational thresholds such as pressure spikes, maintenance events including vibration reductions following component replacement, and state transitions characterized by pressure shifts during mode changes. Events are generated via a multivariate normal distribution with correlations ranging from 0.3 to 0.5 to simulate realistic dependencies between sensor streams. Savitzky-Golay filters [24] smooth the generated signals to remove unrealistic high-frequency noise while preserving the characteristic timescales of industrial processes. This design balances realism with computational efficiency, prioritizing rapid prototyping while features like operational hours and system state enhance contextual modeling for RL and symbolic reasoning.
For validation, we conducted statistical tests including Kolmogorov-Smirnov and Anderson-Darling tests [29], supplemented by cross-correlation analysis to ensure data fidelity. The visual representations in Fig. 3 illustrate the characteristics of our synthetic data. Panel (a) shows vibration measurements over operational hours with labeled event types. Panel (b) displays the distribution of vibration sensor values, distinguishing between normal operational data and key events through color coding. Panel (c) demonstrates pressure readings during a specific time window, highlighting dynamic changes in system state. These visualizations confirm that the synthetic dataset captures realistic patterns and event characteristics suitable for evaluating ANSR-DT’s detection and adaptation capabilities.
@@ IV-B1 Deep Learning Component
The CNN-LSTM neural architecture comprises two Conv1D layers with 64 and 128 filters (kernel size 3), followed by batch normalization and max-pooling. Two LSTM layers with 100 and 50 units then process the spatially-extracted features, with an attention mechanism prioritizing critical time steps within the temporal sequences. This configuration totals 189,481 parameters, carefully balanced to provide sufficient capacity for the 5,000-sample dataset without overfitting on the relatively compact training set.
For optimization, we employed Bayesian optimization [30] to tune hyperparameters including learning rates spanning $10^{-4}$ to $10^{-2}$, dropout rates from 0.1 to 0.5, and network architecture configurations. The final configuration used a learning rate of $10^{-3}$ and dropout rate of 0.3 based on validation performance across the hyperparameter search space. Time series cross-validation with a walk-forward approach ensured robust performance on sequential data, where each fold respects temporal ordering to prevent information leakage from future to past. Early stopping with patience of 10 epochs prevented overfitting by terminating training when validation loss ceased improving. The model was trained using the Adam optimizer with batch size 64 for 20 epochs, achieving convergence as shown in Fig. 5.
@@ IV-B2 Symbolic Reasoning Component
The symbolic component translates neural network outputs into interpretable logical rules and facts using a Prolog-compatible format [27]. The knowledge base contains domain-specific predicates including sensor_reading(Sensor, Value, Time) to encode raw measurements, system_state(State, Time) to represent operational modes, and key_event(Type, Confidence, Time) to flag detected anomalies. These predicates enable logical inference over temporal sequences of observations, supporting queries about system state evolution and event causality.
The framework extracted 14 rules from test predictions, with the rule count stabilizing at 14 across PPO training durations of 10,240 and 200,704 timesteps. This consistency, despite varying RL policies, demonstrates the symbolic component’s reliability in translating neural predictions into interpretable rules on the condensed dataset. Neural rule 1 flags operational deviations with confidence 0.999 when all metrics—temperature, vibration, pressure, efficiency index—fall below their respective thresholds. This rule activated during the inference scenario with pressure reading 18.0, demonstrating actionable symbolic reasoning. Neural rule 6 combines gradient-based conditions (rising efficiency index, temperature, and vibration) with threshold-based conditions, achieving confidence 0.865 but zero test set activations, suggesting it captures rare but potentially important patterns.
@@
For each candidate rule, we calculated confidence scores using the formula shown in Eq. 4, where $\text{support}(r)$ is the frequency of rule $r$’s condition being satisfied, $\text{support}(r\Rightarrow c)$ is the frequency of both the condition and conclusion being satisfied, and $\text{TP}/(\text{TP}+\text{FP})$ is the precision of the rule’s predictions. Rules with confidence scores above threshold $\tau=0.85$ were incorporated into the knowledge base. Conflicting rules were resolved through confidence-based prioritization, where higher-confidence rules take precedence, and specificity analysis, where more specific rules (those with more conditions) override more general rules when both apply. Most learned rules achieved confidence scores above 0.9, demonstrating high reliability in rule extraction.
@@ IV-B3 Neuro-Symbolic Integration Process
The integration between neural and symbolic components operates through a bidirectional interface ensuring seamless information flow [18]. Neural network outputs are converted into symbolic facts using a thresholding mechanism, where predictions exceeding a specified confidence threshold become symbolic facts available for reasoning. Patterns detected by the neural component are analyzed to extract potential rules, which are then evaluated for confidence against the observed data. Rules exceeding the confidence threshold $\tau$ are added to the knowledge base, subject to conflict resolution as described above.
Symbolic insights remained identical across PPO training runs of 10,240 and 200,704 timesteps, consistently including System Stress Triggered and activations of neural_rule_1 and neural_rule_4. This stability, despite a refined RL policy, enhances interpretability by ensuring reliable, repeatable explanations for human operators—a key advantage in safety-critical settings where decision rationales must be consistent and auditable. Table I summarizes the input and output specifications for each major component, illustrating how data flows through the integrated system.
@@ IV-C Reinforcement Learning Implementation
ANSR-DT implements the PPO algorithm [10] for adaptive decision-making in a custom OpenAI Gym-compatible environment. The environment encapsulates sensor readings, system states, and derived metrics as observation space, while providing operational adjustments as action space. Each state combines raw sensor readings (temperature, vibration, pressure), processed features (operational hours, efficiency index), and symbolic facts (activated rules, current system state) to provide a comprehensive view of the system. The action space includes continuous control parameters such as operational settings and threshold adjustments, and discrete decisions such as maintenance scheduling and mode switching.
The multi-objective reward function balances efficiency, user satisfaction, and safety as formalized in Eq. 5. Efficiency measures how well the system maintains performance metrics near target values. Satisfaction captures alignment with operator preferences and historical best practices. Safety penalizes actions that could lead to hazardous conditions or violate operational constraints. The weighting coefficients $\alpha_{1}=0.5$, $\alpha_{2}=0.3$, and $\alpha_{3}=0.2$ reflect the relative importance of each objective, prioritizing efficiency while ensuring that safety and user preferences are adequately considered.
@@
The policy network is implemented as a multi-layer perceptron with three hidden layers containing 256, 128, and 64 units respectively, using ReLU activations. This architecture maps states to action probabilities for discrete actions and action means for continuous actions. The training configuration employed a clipping parameter $\epsilon=0.2$ to constrain policy update magnitude, a discount factor $\gamma=0.99$ for balancing immediate and future rewards, a GAE parameter $\lambda=0.95$ for advantage estimation, and a learning rate of $10^{-3}$ with 10 epochs per policy update and batch size 64.
@@ IV-D Dynamic Rule Updating Mechanism
The continuous rule updating mechanism adapts the symbolic knowledge base based on reinforcement learning insights. Rule candidate generation occurs when the neural component detects novel patterns or when RL agent experiences suggest that existing rules may be incomplete or suboptimal. Each candidate rule’s confidence is assessed based on its predictive accuracy (how often it correctly identifies key events) and support in the observed data (how frequently its conditions are satisfied). Rules exceeding the confidence threshold are incorporated into the knowledge base according to Eq. 6, where existing rules are augmented with new facts derived from recent observations and learning outcomes.
@@
Conflicting rules are resolved through confidence-based prioritization, where higher-confidence rules take precedence when multiple rules apply to the same situation. Specificity analysis ensures that more detailed rules (those with more conditions) override more general rules to capture nuanced operational contexts. A periodic rule pruning process removes rules with decreasing confidence or relevance, identified through tracking of activation frequency and predictive accuracy over sliding time windows. This dynamic updating ensures that the symbolic reasoning component evolves alongside changing operational conditions and learning insights.
IV-E Application Scenarios
To demonstrate ANSR-DT’s practical utility, we implemented three industrial application scenarios derived from our synthetic data. The predictive maintenance scenario monitors equipment health through temperature and vibration sensors to predict maintenance needs before critical failures occur. The neural component identifies vibration signatures indicative of bearing wear, manifested as increased amplitude in specific frequency bands. Symbolic reasoning provides interpretable maintenance recommendations using rules such as high_frequency_vibration(X) $\land$ rising_temperature(X) $\rightarrow$ maintenance_required(X, urgent). The system learns optimal maintenance scheduling based on operational constraints (minimizing downtime) and historical outcomes (balancing preventive maintenance costs against failure costs).
The process optimization scenario focuses on optimizing operational parameters for efficiency. This implementation integrates temperature, pressure, and derived efficiency metrics to monitor process performance continuously. The RL component learns optimal operational settings under varying conditions, discovering that certain temperature-pressure combinations maximize efficiency for different product types or environmental conditions. Symbolic rules explain why specific settings are optimal in given contexts, enabling operators to understand and validate the system’s recommendations. The system adapts to operator preferences while maintaining safety constraints, ensuring that optimization never compromises operational safety margins.
The deviation detection and response scenario demonstrates ANSR-DT’s capabilities in identifying and responding to operational anomalies. This implementation uses multi-sensor fusion to analyze temperature, vibration, and pressure data for complex event patterns that may not be apparent from any single sensor stream. The system provides graduated responses prioritized by event severity and confidence, ranging from flagging low-confidence anomalies for operator review to automatically triggering protective actions for high-confidence critical events. Continuous learning allows adaptation to new deviation patterns, with symbolic reasoning offering operators clear explanations of detected operational deviations and recommended responses [31].
IV-F Integration and System Architecture
We integrated all components into a cohesive pipeline with well-defined interfaces between layers, as illustrated in Fig. 4. The implementation includes a data preprocessing pipeline for sensor data normalization and fusion, handling missing values through interpolation and detecting measurement outliers through statistical process control. The neural component is implemented in TensorFlow 2.8, leveraging GPU acceleration for training and inference. A symbolic reasoning engine using ProbLog 2.2 [27] handles probabilistic logic inference over the extracted rule base. The reinforcement learning module, based on Stable Baselines3 [32], provides robust implementations of PPO with vectorized environments for efficient training. These components connect through a Python-based middleware with standardized interfaces defined through abstract base classes, ensuring modularity and extensibility. Visualization tools monitor system performance in real-time and explain decisions through interactive knowledge graphs and rule activation traces, as shown in Fig. 8.
@@ IV-G Future Extensions
Based on our implementation experience, several promising directions for future enhancements emerge. Scaling to larger datasets will test rule management beyond the current 14 rules, requiring development of hierarchical rule structures to organize large rule bases. Diversifying rule extraction to capture more complex patterns, such as temporal sequences of events or conditional dependencies between sensor streams, will leverage extended PPO training for enhanced adaptability. Adapting the framework for deployment on edge devices with constrained computational resources will require model compression and quantization techniques. Integrating natural language processing will enable more intuitive human-machine interaction, allowing operators to query the system in natural language and receive explanations in accessible terms. Evaluating the framework’s applicability in diverse domains beyond manufacturing, such as energy systems or transportation networks, will assess generalization capabilities [33]. These extensions aim to further enhance the adaptability, interpretability, and practical utility of ANSR-DT for complex industrial applications.
@@ V-A Experimental Setup
Experiments were conducted using synthetic industrial data with a 60/20/20 train/validation/test split containing 5% labeled dynamic patterns distributed across operational thresholds, maintenance patterns, state transitions, and interaction deviations. To address class imbalance in the dataset, we applied class weighting (0: 64.7, 1: 0.5) during model training, ensuring that the minority class (key events) received appropriate emphasis during gradient computation. We established a rigorous comparison framework against CNN-LSTM baseline models to isolate the contributions of the neuro-symbolic integration and reinforcement learning components.
V-B Performance Analysis and Time-Series Pattern Extraction
Table II presents comparative performance metrics between ANSR-DT and the CNN-LSTM baseline. Our framework achieves a peak validation accuracy of 99.5% by epoch 15 across 20 training epochs, compared to 86.87% for the CNN-LSTM baseline. However, fluctuations observed during training (e.g., validation accuracy dipping to 55% at epoch 16) suggest potential overfitting on the reduced dataset, mitigated through early stopping and class weighting to address imbalance. The training progression illustrated in Fig. 5 shows convergence with validation accuracy reaching its peak at epoch 15, indicating efficient learning despite some fluctuations attributable to the smaller dataset size.
@@ V-B Performance Analysis and Time-Series Pattern Extraction
Accuracy (%) 99.5 (peak) 86.87 +12.63 << 0.01
Precision 0.80 $\pm$ 0.05 0.9358 -14.50 << 0.05
Recall 0.75 $\pm$ 0.04 0.6355 +17.82 << 0.05
@@
The most substantial improvement appears in recall performance, with a 17.82% increase ($p<0.05$) over the baseline, indicating significantly enhanced capability to identify critical patterns with fewer false negatives—a crucial characteristic for industrial safety applications where missing operational deviations can lead to equipment damage or safety hazards. This high detection accuracy is achieved through the combination of CNN-LSTM architecture with an attention mechanism, which effectively learns temporal dependencies across multiple sensor streams. The precision-recall trade-off, deliberately engineered to prioritize recall in safety-critical industrial settings, reflects the operational reality that false negatives typically carry higher operational risk than false positives. While precision decreased by 14.50%, this trade-off is acceptable given that false alarms can be reviewed by operators, whereas missed critical events may propagate into serious failures.
Fig. 6 provides visual confirmation of classification performance across decision thresholds. The confusion matrix in panel (a) shows high true positive and true negative rates, with 343 true negatives (100.0%) and 153 true positives (98.1%), minimal false negatives (3 samples, 1.9%), and no false positives. This distribution confirms the model’s ability to maintain high specificity while achieving strong sensitivity. The precision-recall curve in panel (b) demonstrates robust performance across different decision thresholds with an Average Precision score of 1.00, indicating that the model maintains high precision even as recall increases—a desirable characteristic for operational deployment where threshold selection must balance detection sensitivity with alarm rate.
[…]